# Representation and Integration by Parts Formulas for Affine Processes

### v1

@@ 3 Derivatives of expectations of affine diffusions
Here, $\mathrm{Log}$ denotes the principal value of the complex logarithm, that is, the imaginary part of $\mathop{\mathrm{Log}}z$ lies in the interval $(-\pi,\pi]$. That is,
$E[\exp(i\theta X^{x}_{t})]=\begin{dcases}\biggl(\frac{2}{2-\alpha i\theta}\biggr)^{2b/\alpha}\exp\biggl(\frac{2i\theta}{2-\alpha i\theta}x\biggr);&\text{ if }\beta=0,\\ \biggl(\frac{2\beta}{2\beta-(e^{\beta t}-1)\alpha i\theta}\biggr)^{2b/\alpha}\exp\biggl(\frac{2\beta e^{\beta t}i\theta}{2\beta-(e^{\beta t}-1)\alpha i\theta}x\biggr);&\text{ if }\beta\neq 0.\end{dcases}$
Note that the limit of $E[\exp(i\theta X^{x}_{t})]$ as $\beta\to 0$ coincides with the case $\beta=0$. That is, $\beta\mapsto E[\exp(i\theta X^{x}_{t})]$ is continuous at 0.
@@ 3.1 Derivative with respect to the initial value
First, we denote the characteristic function of $X^{x}_{t}$ (resp. $X^{1,x}_{t}$) by $\varphi^{x}_{t}$ (resp. $\varphi^{1,x}_{t}$). Moreover, we denote the law of $X^{x}_{t}$ (resp. $X^{1,x}_{t}$) by $\nu^{x}_{t}$ (resp. $\nu^{1,x}_{t}$). Then, by using inverse Fourier transform,
@@
Here, we have applied Fubini’s theorem as $(\theta,y)\mapsto\widehat{f}(\theta)e^{i\theta y}$ is $(\mathrm{Leb}\times\nu^{x}_{t})$-integrable by assumption, where $\mathrm{Leb}$ denotes the Lebesgue measure. Moreover, we have
$|\partial_{x}\varphi^{x}_{t}(\theta)|$ $\leq\biggl|\frac{2}{2-\alpha i\theta}\biggr||\varphi^{x}_{t}(\theta)|$
$=\frac{2}{2+\alpha^{2}\theta^{2}}|\varphi^{x}_{t}(\theta)|$
$\leq\frac{2}{2}\cdot 1$
$=1.$
Thus,
$|\widehat{f}(\theta)\,\partial_{x}\varphi^{x}_{t}(\theta)|\leq|\widehat{f}(\theta)|,$
@@
$E[f(X^{1,x}_{t})]=\frac{1}{2\pi}\int_{\mathbb{R}}\widehat{f}(\theta)\varphi^{1,x}_{t}(\theta)\,dy\,d\theta.$
Now, we first assume that $\beta=0$. In this case,
$\varphi^{x}_{t}(\theta)=\biggl(\frac{2}{2-\alpha i\theta}\biggr)^{2b/\alpha}\exp\biggl(\frac{2i\theta}{2-\alpha i\theta}x\biggr).$
Therefore,
$\partial_{x}\varphi^{x}_{t}(\theta)=\biggl(\frac{2}{2-\alpha i\theta}\biggr)^{2b/\alpha+1}i\theta\exp\biggl(\frac{2i\theta}{2-\alpha i\theta}x\biggr).$
@@
$\partial_{x}\varphi^{x}_{t}(\theta)$ $=\frac{2}{\alpha t}\biggl(\biggl(\frac{2}{2-\alpha i\theta}\biggr)^{2b/\alpha+1}\exp\biggl(\frac{2i\theta}{2-\alpha i\theta}x\biggr)-\biggl(\frac{2}{2-\alpha i\theta}\biggr)^{2b/\alpha}\exp\biggl(\frac{2i\theta}{2-\alpha i\theta}x\biggr)\biggr)$
@@
$=\int_{\mathbb{R}}\frac{1}{2\pi}\int_{\mathbb{R}}\widehat{f^{\prime}}(\theta)e^{i\theta x}\,d\theta\,\nu^{x}_{t}(dy)$
$=\frac{1}{2\pi}\int_{\mathbb{R}}\int_{\mathbb{R}}\widehat{f^{\prime}}(\theta)e^{i\theta x}\,\nu^{x}_{t}(dy)\,d\theta$
@@
IBP (Thm. 2) $b\leadsto b-\alpha/2$ $E[f^{\prime}(X^{x}_{t})]=C_{2}(E[f(X^{-1,x}_{t})]-E[f(X^{x}_{t})])$
@@ 4 An extension to particularly simple multidimensional case
$\partial_{x_{k}}E[f(X^{x}_{t})]=\begin{dcases}\frac{2}{\alpha_{k}t}\bigl(E[f(X^{e_{k},x}_{t})]-E[f(X^{x}_{t})]\bigr);&\text{ if }\beta_{k,k}=0,\\ \frac{2\beta_{k,k}}{\alpha_{k}(e^{\beta_{k,k}t}-1)}\bigl(E[f(X^{e_{k},x}_{t})]-E[f(X^{x}_{t})]\bigr);&\text{ if }\beta_{k,k}\neq 0.\end{dcases}$
@@
$=\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\widehat{f}(\theta)e^{i\langle\theta,y\rangle}d\theta\,\nu_{X^{x}_{t}}(dy)$
$=\int_{\mathbb{R}^{d}}\widehat{f}(\theta)\int_{\mathbb{R}^{d}}e^{i\langle\theta,y\rangle}\nu_{X^{x}_{t}}(dy)\,d\theta\$
$=\int_{\mathbb{R}^{d}}\widehat{f}(\theta)\varphi_{X^{x}_{t}}(\theta)\,d\theta,$
@@ 5.1 CIR model
$E[f^{\prime}(r_{t})]=\frac{-2ke^{-kt}}{\sigma^{2}(e^{-kt}-1)}\bigl(E[f(r_{t})]-E[f(r^{\langle-1\rangle}_{t})]\bigr).$
@@ 5.2 Model of evolution of population
where $x$ is a positive integer and $(\xi_{n,k})_{n,k}$ are IID random variables which model the descendants of the $k$-th member of the $n$-th degeneration.

### v2

@@ 3 Derivatives of expectations of affine diffusions
Here, $\mathrm{Log}$ denotes the principal value of the complex logarithm, that is, the imaginary part of $\mathop{\mathrm{Log}}z$ lies in the interval $(-\pi,\pi]$.
Thus, the characteristic function of $X^{x}_{t}$ is given as follows.
$E[\exp(i\theta X^{x}_{t})]=\begin{dcases}\biggl(\frac{2}{2-\alpha i\theta t}\biggr)^{2b/\alpha}\exp\biggl(\frac{2i\theta}{2-\alpha i\theta t}x\biggr);&\text{ if }\beta=0,\\ \biggl(\frac{2\beta}{2\beta-(e^{\beta t}-1)\alpha i\theta}\biggr)^{2b/\alpha}\exp\biggl(\frac{2\beta e^{\beta t}i\theta}{2\beta-(e^{\beta t}-1)\alpha i\theta}x\biggr);&\text{ if }\beta\neq 0.\end{dcases}$
Hereafter, throughout this paper, for a complex number $\alpha$, its real power $\alpha^{c}$ is understood to represent $e^{c\mathop{\mathrm{Log}}\alpha}$. Note that the limit of $E[\exp(i\theta X^{x}_{t})]$ as $\beta\to 0$ coincides with the case $\beta=0$. That is, $\beta\mapsto E[\exp(i\theta X^{x}_{t})]$ is continuous at 0.
@@ 3.1 Derivative with respect to the initial value
In the following, we denote the characteristic function of $X^{x}_{t}$ (resp. $X^{1,x}_{t}$) by $\varphi^{x}_{t}$ (resp. $\varphi^{1,x}_{t}$). Moreover, we denote the law of $X^{x}_{t}$ (resp. $X^{1,x}_{t}$) by $\nu^{x}_{t}$ (resp. $\nu^{1,x}_{t}$). Then, by using inverse Fourier transform,
@@
Here, we have applied Fubini’s theorem as $(\theta,y)\mapsto\widehat{f}(\theta)e^{i\theta y}$ is $(\mathrm{Leb}\times\nu^{x}_{t})$-integrable by assumption, where $\mathrm{Leb}$ denotes the Lebesgue measure.
First, we consider the case $\beta=0$. In this case,
$\varphi^{x}_{t}(\theta)=\biggl(\frac{2}{2-\alpha i\theta t}\biggr)^{2b/\alpha}\exp\biggl(\frac{2i\theta}{2-\alpha i\theta t}x\biggr).$
Therefore,
$\partial_{x}\varphi^{x}_{t}(\theta)=\biggl(\frac{2}{2-\alpha i\theta t}\biggr)^{2b/\alpha+1}i\theta\exp\biggl(\frac{2i\theta}{2-\alpha i\theta t}x\biggr).$
Moreover, we have
$|\partial_{x}\varphi^{x}_{t}(\theta)|$ $\leq\biggl|\frac{2i\theta}{2-\alpha i\theta t}\biggr||\varphi^{x}_{t}(\theta)|$
$=\frac{2|\theta|}{\sqrt{4+\alpha^{2}\theta^{2}t^{2}}}|\varphi^{x}_{t}(\theta)|$
$\leq\frac{2|\theta|}{\alpha t|\theta|}\cdot 1$
$=\frac{2}{\alpha t}.$
for $\theta\neq 0$. Thus,
$|\widehat{f}(\theta)\,\partial_{x}\varphi^{x}_{t}(\theta)|\leq\frac{2}{\alpha t}|\widehat{f}(\theta)|,$
@@
$E[f(X^{1,x}_{t})]=\frac{1}{2\pi}\int_{\mathbb{R}}\widehat{f}(\theta)\varphi^{1,x}_{t}(\theta)\,d\theta.$
@@
$\partial_{x}\varphi^{x}_{t}(\theta)$ $=\biggl(\frac{2}{2-\alpha i\theta t}\biggr)^{2b/\alpha+1}i\theta\exp\biggl(\frac{2i\theta}{2-\alpha i\theta t}x\biggr)$
$=\biggl(\frac{2}{2-\alpha i\theta t}\biggr)^{2b/\alpha+1}\frac{2}{\alpha t}\biggl(1-\frac{2-\alpha i\theta t}{2}\biggr)\exp\biggl(\frac{2i\theta}{2-\alpha i\theta t}x\biggr)$
$=\frac{2}{\alpha t}\biggl(\biggl(\frac{2}{2-\alpha i\theta t}\biggr)^{2b/\alpha+1}\exp\biggl(\frac{2i\theta}{2-\alpha i\theta t}x\biggr)-\biggl(\frac{2}{2-\alpha i\theta t}\biggr)^{2b/\alpha}\exp\biggl(\frac{2i\theta}{2-\alpha i\theta t}x\biggr)\biggr)$
@@
$=\int_{\mathbb{R}}\frac{1}{2\pi}\int_{\mathbb{R}}\widehat{f^{\prime}}(\theta)e^{i\theta y}\,d\theta\,\nu^{x}_{t}(dy)$
$=\frac{1}{2\pi}\int_{\mathbb{R}}\int_{\mathbb{R}}\widehat{f^{\prime}}(\theta)e^{i\theta y}\,\nu^{x}_{t}(dy)\,d\theta$
@@
IBP (Thm. 2) $b\leadsto b-\alpha/2$ $E[f^{\prime}(X^{x}_{t})]=C_{2}(E[f(X^{x}_{t})]-E[f(X^{-1,x}_{t})])$
@@ 4 An extension to particularly simple multidimensional case
$\partial_{x_{k}}E[f(X^{x}_{t})]=\begin{dcases}\frac{2}{\alpha_{k}t}\bigl(E[f(X^{e_{k},x}_{t})]-E[f(X^{x}_{t})]\bigr);&\text{ if }\beta_{k,k}=0,\\ \frac{2\beta_{k,k}e^{\beta_{k,k}t}}{\alpha_{k}(e^{\beta_{k,k}t}-1)}\bigl(E[f(X^{e_{k},x}_{t})]-E[f(X^{x}_{t})]\bigr);&\text{ if }\beta_{k,k}\neq 0.\end{dcases}$
@@
$=\int_{\mathbb{R}^{d}}\frac{1}{(2\pi)^{d}}\int_{\mathbb{R}^{d}}\widehat{f}(\theta)e^{i\langle\theta,y\rangle}d\theta\,\nu_{X^{x}_{t}}(dy)$
$=\frac{1}{(2\pi)^{d}}\int_{\mathbb{R}^{d}}\widehat{f}(\theta)\int_{\mathbb{R}^{d}}e^{i\langle\theta,y\rangle}\nu_{X^{x}_{t}}(dy)\,d\theta\$
$=\frac{1}{(2\pi)^{d}}\int_{\mathbb{R}^{d}}\widehat{f}(\theta)\varphi_{X^{x}_{t}}(\theta)\,d\theta,$
@@ 5.1 CIR model
$E[f^{\prime}(r_{t})]=\frac{-2k}{\sigma^{2}(e^{-kt}-1)}\bigl(E[f(r_{t})]-E[f(r^{\langle-1\rangle}_{t})]\bigr).$
@@ 5.2 Model of evolution of population
where $x$ is a positive integer and $(\xi_{n,k})_{n,k}$ are IID random variables which model the descendants of the $k$-th member of the $n$-th generation.